### v1

# Polaritonic control of blackbody infrared radiative dissociation

Abstract
Vibrational strong light-matter coupling offers a promising approach for controlling chemical reactivity with infrared microcavities. This study explores the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes. Using a Master equation approach, we simulate the effects of infrared field confinement and vibrational strong coupling on BIRD rates for diatomic molecules. We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process. Our findings reveal conditions for significant enhancement and mild suppression of radiative dissociation, establishing upper bounds for BIRD rates under weak and strong coupling. These results provide new strategies and limitations for controlling reactive processes with infrared resonators.
I Introduction
Chemical reaction control is critical in applications ranging from industrial catalysis to medicine. Conventional methods for manipulating reaction rates involve varying temperature, pressure, or catalyst composition. An unconventional approach for controlling thermal reaction kinetics has been recently reported Thomas et al. (2016, 2019); Hirai et al. (2020); Ahn et al. (2023) employing infrared (IR) microcavities Kavokin et al. (2017) consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations Nagarajan, Thomas, and Ebbesen (2021). When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device Vinogradov (1992); Shalabney et al. (2015); Simpkins et al. (2015). Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction ratesGalego et al. (2019); Campos-Gonzalez-Angulo, Ribeiro, and Yuen-Zhou (2019); Zhdanov (2020); Wang, Flick, and Yelin (2022); Lindoy, Mandal, and Reichman (2022); Schäfer et al. (2022); Sun and Vendrell (2022); Ahn et al. (2023); Montillo Vega, Ying, and Huo (2024); Moiseyev (2024). Still, several fundamental questions remain Simpkins, Dunkelberger, and Owrutsky (2021); Campos-Gonzalez-Angulo et al. (2023), including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted. In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation Dunbar (2004) under weak and strong coupling with a suitable material.
Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation Perrin (1970). However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism Lindemann et al. (1922); Daniels (1928); King and Laidler (1984). Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressuresThoelmann, Tonner, and McMahon (1994); Dunbar et al. (1995); Schnier et al. (1996). Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers Dunbar (2004).
BIRD rates are heavily influenced by the thermal radiation spectrum. Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different Kavokin et al. (2017), microcavity BIRD is expected to proceed with different rates compared to free space BIRD. However, this difference’s order of magnitude, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime. Our study provides answers to these questions. Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but our qualitative conclusions are generalizable to polyatomic systems.
The primary aim of this study is to characterize thermal radiative dissociation in microcavities. We employ a Master equation treatment to investigate blackbody infrared radiative dissociation for diatomic molecules in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material. Our results provide upper bounds for enhancement and suppression of BIRD in weakly and strongly coupled microcavities and identify overtone transitions as key ingredients for the observed modulation of BIRD. In Sec. II, we describe our methodology, Sec. III provides our main results and related discussion, and Sec. IV summarizes our main results and conclusions.
II Methods
Figure 1: (a) Illustration of the Morse potential describing the vibrational states and transitions of the HF molecule. (b) and (c): A scheme illustrating an HF molecule inside a planar cavity and its dispersion under the weak coupling regime, respectively. (d) and (e): A scheme illustrating an HF molecule inside a planar cavity with the host material strongly coupled to the cavity and the dispersion of polariton states, respectively.
II.1 Master Equation for diatomic BIRD
Our analysis of BIRD in microcavities employs the same tools successfully applied to investigate this process in free space Dunbar (2004); Dunbar et al. (1995); Stevens et al. (2002); Dunbar and McMahon (1998). These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold. Rotational degrees of freedom are ignored since at the temperatures of interest to BIRD, the rotational dynamics is essentially classical, only introducing a nonzero linewidth to the vibrational absorption and emission spectrum. We discuss further implications of this condition in the context of our results below.
The rate of change of the population of the $i$th vibrational state ($N_{i}$) is given by
$\frac{dN_{i}(t)}{dt}=\sum_{j\neq i}^{i_{\text{max}}}\left[k_{ij}N_{j}(t)-k_{ji}N_{i}(t)\right]-k_{\text{loss}}N_{i}(t)\delta_{i,i_{\text{max}}}\;,$ (1)
where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ refers to the highest energy bound state (Fig. 1a). We assume that the state $i_{\text{max}}$ undergoes irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$. For the sake of simplicity, we have assumed the thermalization of the background radiation is faster than all considered radiative transitions. Equation 1 can be written in matrix notation as
$\frac{d[\textbf{N}]}{dt}=-\textbf{J}\cdot[\textbf{N}],$ (2)
where N is a column vector containing the vibrational populations and $\mathbf{J}$ is the transition matrix. The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction Valance and Schlag (1966); Gilbert and Smith (1990).
II.2 Radiative transition rates
Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule Fermi (1932); Sakurai and Napolitano (2017)) and written in terms of Einstein coefficients in free space Einstein (1917); Cohen-Tannoudji, Diu, and Laloë (2019) and their generalizations in a weakly coupled and polaritonic microcavity. Rotational dynamics is much faster than thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
In free space, the spontaneous emission rate is given by the Einstein $A$ coefficientEinstein (1917); Cohen-Tannoudji, Diu, and Laloë (2019). In the case where spontaneous emission induces the $i\rightarrow j$ transition with $i>j$ this coefficient is given by
$k_{ij}^{\text{SpEm}}=A_{ij}=\frac{\omega_{ij}|\mu_{ij}|^{2}\pi}{3\epsilon_{0}\hbar}D_{0}(\omega_{ij}),$ (3)
where $\epsilon_{0}$ and $\hbar$ are the vacuum electrical permittivity and reduced Plank’s constant, respectively, $\omega_{ij}=(E_{i}-E_{j})/\hbar$ is the transition frequency, $\mu_{ij}$ is the corresponding transition dipole moment and $D_{0}(\omega)=\omega^{2}/\pi^{2}c^{3}$ is the free-space radiative density of states (DOS). The Einstein $B$ coefficient associated with stimulated emission and absorption processes involving states $i$ and $j$ is given by $B_{ij}=\pi|\mu_{ij}|^{2}/(3\epsilon_{0}\hbar^{2})$. The corresponding thermal stimulated emission and absorption rate constants are given by
$k_{ij}^{\text{Abs}}(T)=k_{ij}^{\text{StEm}}(T)=B_{ij}\rho_{0}(\omega_{ij},T)\;,$ (4)
where $\rho_{0}(\omega,T)$ is the free space blackbody radiation energy density
$\rho_{0}(\omega,T)=\hbar\omega~{}n_{\text{BE}}(\omega,T)D_{0}(\omega),$ (5)
where $n_{\text{BE}}(\omega,T)$ is the Bose-Einstein distribution thermal occupation number $n_{\text{BE}}(\omega,T)=1/[\text{exp}(\hbar\omega/k_{B}T)-1]$Bose (1924) and $k_{B}$ is the Boltzmann constant.
The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$. We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic molecules is given by the Morse potential [Fig. 1(a)]Morse (1929). The bound wave functions associated with the Morse potential are also employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
$\mu_{ij}$ $=\int\psi_{i}(r)\mu(r)\psi_{j}(r)\mathrm{d}r,$ (6)
where $\psi_{i}(r)$ represents the $i$th vibrational level wave function, and the dipole function $\mu(r)$ models the nonlinear change in the (electronic ground-state) molecular electrical dipole with the bond length $r$ Bishop (1990).
We investigated the IR radiative dissociation of hydrogen fluoride, HF, and lithium hydride, LiH. The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-1 were obtained from previous work on laser-induced HF dissociationKaluža et al. (1994). The LiH dipole function was interpolated from ab initio quantum chemical computations (see Supplementary Information). Morse potential parameters include the equilibrium bond length ($r_{e}$), electronic ground-state dissociation energy ($D_{e}$), and harmonic frequency ($\omega_{e}$) of each molecule and are listed in Table 1.
Table 1: Parameters used for Morse potential of HF Stine and Noid (1979); Guldberg and Billing (1991); Gross, Neuhauser, and Rabitz (1993); Kaluža et al. (1994) and LiH Irikura (2007).
Constant HF LiH
$D_{e}$ (eV) 6.123 2.641
$r_{e}$ (Å) 0.926 1.595
$\nu_{e}$ (cm-1) 4138 1405
$\nu_{e}\chi_{e}$ (cm-1) 86.7 28.3
$m_{r}$ (a.u) 0.95 0.88
Note that the examined diatomic species have severely reduced room temperature BIRD rates relative to typical experiments involving polyatomic ions Dunbar and McMahon (1998); Dunbar et al. (1995) due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation. This feature motivated our choice of examining diatomic BIRD at temperatures $T=\mathcal{O}(\hbar\omega_{e}/k_{\text{B}})$, an order of magnitude greater than room temperature. Such temperatures lead to numerically stable BIRD rates in the various regimes considered in this work. Despite their largely different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior. This independence to microscopic details demonstrates the robustness of our analysis.
The transition rates $k_{ij}$ are proportional to the electromagnetic density of states $D(\omega)$, suggesting BIRD rates can be controlled by placing the reactive molecule inside a microcavity. To examine this scenario, we adapt the free space formalism described above to arbitrary photonic devices by replacing $D_{0}(\omega)$ in Eqs. 3 and 5 with the appropriate DOS for a weakly coupled microcavity $D_{\text{C}}(\omega)$.
In a planar empty Fabry-Perot microcavity composed of two parallel mirrors (with unit reflectivity for simplicity) separated by a distance $L_{C}$, [Fig. 1(b)], the frequencies of the electromagnetic modes are given by
$\omega_{\text{C}}(m,\textbf{q})=c\sqrt{q^{2}+\frac{m^{2}\pi^{2}}{L_{C}^{2}}}\;,$ (7)
where $\mathbf{q}=(q_{x},q_{y})\in\mathbb{R}^{2}$ is the in-plane projection of the mode wave vector, $q^{2}=q_{x}^{2}+q_{y}^{2}$, and $m$ is an integer quantum number that identifies the longitudinal wave vector $k_{z}=m\pi/L_{C}$ and thus the photonic energy bands in Fig. 1(c). For transverse-magnetic (TM) modes $m=0,1,2,...$ whereas $m=1,2,3...$ for transverse-electric (TE) modes Zoubi and La Rocca (2005); Jackson (2021). Including the TM0 mode here is crucial, as otherwise, transitions with energy below the cavity cutoff would be impossible.
Using Eq. 7, it can be shown (see Supporting Information) that the spatially-averaged electromagnetic DOS inside a microcavity in the weak coupling regime can be expressed by Barnes, Horsley, and Vos (2020)
$D_{C}(\omega)=\frac{\omega}{\pi c^{2}L_{C}}\left\lfloor\frac{\omega L_{C}}{\pi c}\right\rfloor+\frac{\omega}{2\pi c^{2}L_{C}}\;,$ (8)
where $\lfloor x\rfloor$ denotes the floor function.
Lastly, we considered the BIRD scenario in which the diatomic interacts weakly with the electromagnetic component of polariton modes emergent from vibrational strong coupling between the microcavity modes and a material system (solid-state or molecular ensemble) with sufficiently large IR oscillator strength. We emphasize this material system does not undergo BIRD; its role is to generate polariton normal-modes with a tunable density of states and thermal energy density, as shown in Figs. 1(d) and (e). As we show in the SI, the formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
The Hamiltonian of the strongly coupled molecular subsystem and the microcavity is given in the Power-Zienau-Woolley (dipole) gauge Power and Zienau (1959); Woolley (1971)
$H=H_{\text{SM}}+H_{\text{L}}+H_{\text{INT}}\;,$ (9)
where the interaction $H_{\text{INT}}$ is given by
$H_{\text{INT}}=\frac{1}{\epsilon_{0}}\int\left[-\textbf{D}(\textbf{r})\cdot\textbf{P}(\textbf{r})+\frac{1}{2}\textbf{P}^{2}(\textbf{r})\right]d^{3}\textbf{r},$ (10)
[…]
$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-m\pi c/L_{C}],$ (11)
where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$ (see SI)
$P_{C}(\omega)=\frac{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}}{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}+\omega^{2}_{C}(\omega)\Omega_{R}^{2}},$ (12)
$\omega_{M}$ is the matter oscillator frequency, $\Omega_{R}$ is the collective light-matter interaction strength, and the polariton group velocity ($v_{g}$) is given by
$v_{g}(\omega,q)=\frac{qc^{2}}{2\omega}\left(1\pm\frac{\omega_{C}^{2}-\omega_{M}^{2}+\Omega_{R}^{2}}{\sqrt{\left(\omega_{C}^{2}+\omega_{M}^{2}+\Omega_{R}^{2}\right)^{2}-4\omega_{C}^{2}\omega_{M}^{2}}}\right)\;,$ (13)
where the positive sign applies when $\omega$ belongs to an upper polariton branch, i.e., $\omega>\omega_{C}$, and the negative sign is used if $\omega<\omega_{C}$. In this expression, we have omitted the frequency dependence of $\omega_{C}$ for simplicity.
Our model shows singular behavior for $D_{P}(\omega)$ as $\omega\rightarrow\omega_{M}$. In this limit, both the photon content $P_{C}(\omega)$ and the group velocity $v_{g}(\omega,q)$ approach zero, but the ratio $P_{C}(\omega)/v_{g}(\omega,q)\rightarrow\infty$ resulting in $D_{P}(\omega)$ diverging to positive infinity. This singularity is due to the formation of the well-known stopgap regionLitinskaya and Agranovich (2009); Huang et al. (2003); Gómez Rivas et al. (2004); Todorov and Sirtori (2012) where $D_{P}(\omega)=0$. As previous work shows Satanin et al. (2005), introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$. Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on the type of disorder and coherence loss mechanism. This motivates our choice to work with the formally exact expression given in Eq. 11 and interpret our results as providing upper bounds for the effect of polaritons on BIRD.
III Results and Discussion
III.1 Microcavity-assisted blackbody infrared radiative dissociation
Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity. The tunable electromagnetic environment afforded by the microcavity allows suggests BIRD rates can be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the change induced by the microcavity in the photon DOS.
In Fig. 2, BIRD rates in the weak coupling regime are shown relative to the free space rate ($k_{0}=2.70\times 10^{-6}$ s-1 for HF at $T/\omega_{e}=0.67$ corresponding to $T=4000~{}\text{K}$ and $k_{0}=1.64\times 10^{-5}$ s-1 for LiH at $T/\omega_{e}=0.99$ corresponding to $T=2000~{}\text{K}$). For small microcavity lengths ($L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH), a consistent mild enhancement of the BIRD rate is associated with an overall increase in the microcavity electromagnetic DOS relative to free space over a wide frequency range covering most fundamental transitions and relevant overtones of HF and LiH. The ratios of microcavity BIRD rates to free space BIRD rates quickly approach 1.0 as $L_{C}$ increases beyond a few microns. This is expected since the microcavity DOS converges to the free space limit when $L_{C}\rightarrow\infty$.
The complex oscillatory pattern in the BIRD rates relative to free space shown in Fig. 2 arises from the interplay of transitions enhanced or suppressed depending on $L_{C}$. As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
Figure 2: Ratio of BIRD rate inside a microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for HF and LiH molecules. Red circles highlight specific lengths further discussed in Fig. 3.
To interpret the origin of the oscillatory behavior in Fig. 2 and identify the most consequential radiative transitions involved in BIRD of HF and LiH, we conducted a sensitivity analysis of the dissociation rate. This measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details). For HF dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 23$, with $i\in\{16,17,19\}$ (in the relevant 1500 -- 5000 cm-1 frequency range) are the most consequential for the BIRD rate. Similarly, for LiH, the relevant transitions are $i\rightarrow 29$, with $i\in\{21,22,23,25\}$. In both cases, the identified overtones take the vibrational mode to the highest energy bound state (23 for HF and 29 for LiH) from which dissociation rapidly occurs. We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths.
Figure 3: Microcavity photon density of states normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig. 2. The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color-coding to distinguish between suppression (red) and enhancement (green). Panel (a) shows the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced microcavity photon DOS relative to free space at the highlighted overtones. Panel (b) shows the mild enhancement of HF BIRD in a microcavity with $L_{C}=3.21~{}\mu\text{m}$ can be ascribed to the greater microcavity photon DOS at the frequencies of the highlighted overtones. Panels (c) and (d) show the analogous scenario for LiH, where BIRD is observed to be suppressed at $L_{C}=9.4~{}\mu\text{m}$ but enhanced at $L_{C}=9.6~{}\mu\text{m}$, primarily due to the variation of the microcavity photon density of states at the selected $i\rightarrow 29$ overtone frequencies.
Equipped with the knowledge of the most relevant transitions for BIRD of HF and LiH, we can readily explain the microscopic origin of the intricate pattern in Fig. 2 by considering how the microcavity DOS changes around the previously mentioned overtones. Fig. 3 presents the ratio of the microcavity DOS to the free space for selected values of $L_{C}$ where BIRD suppression or enhancement was verified in Fig. 2 (red circles). Overtone frequencies are highlighted with dotted vertical lines, color-coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity. In Fig. 3a ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig. 2. In contrast, when the microcavity length is slightly increased to $L_{C}=3.21$ $\mu$m (Fig. 3(b) the photon DOS is enhanced at $\omega$ corresponding to the $16\rightarrow 23$ and $19\rightarrow 23$ transitions, and we observe a corresponding leap in the BIRD rate in Fig. 2. Similarly, in Fig. 3(c) ($L_{C}=9.4~{}\mu\text{m})$, all of the relevant LiH overtones but $21\rightarrow 29$ are suppressed, leading to a reduced BIRD rate in this microcavity relative to free space, whereas Fig. 3(d) shows that at $L_{C}=9.6~{}\mu\text{m}$ most of the relevant LiH overtones are enhanced, thus explaining the sudden increase in BIRD rate at $L_{C}=9.6~{}\mu\text{m}$ relative to $L_{C}=9.4~{}\mu\text{m}$.
As we conclude this subsection, we note that suppression effects are nearly negligible in Fig. 2 because the dissociative process has many pathways. Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always nearly negligible as the molecular system explores alternative dissociation pathways. Therefore, our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential. Strictly speaking, we have shown that these conditions rigorously apply to diatomic BIRD, but they are likely to be even more relevant in polyatomic processes.
III.2 Polariton-assisted blackbody infrared radiative dissociation
To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes. Figure 1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photonfrequency-momentum dispersion relation [Fig. 1 (e)] and the electromagnetic DOS. We explore the implications for diatomic BIRD below.
Figure 4 presents the ratio of polariton-assisted HF and LiH BIRD rates to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$. These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 at microcavity length $L_{C}$ fixed at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH, respectively. As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime. We observe prominent peaks for both HF and LiH where the relative dissociation rate is notably enhanced. Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$, where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH. Conversely, as seen in the inset of Figs. 4(a) and (c), when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed. This happens because when $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the relevant overtone transitions fall into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig. 1e). Nonetheless, the suppression is small since, as discussed above, BIRD may proceed via other pathways.
The effect of a variation of collective coupling strengths ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig. 4(b) and (d). When $\omega_{M}-\omega_{i\rightarrow j}$ is around 1-10 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates. However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively. This plateau appears to be insensitive to the magnitude of the Rabi splitting. It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig. 5.
Figure 4: Ratio of polariton-assisted BIRD rates to free space BIRD rates. In the polaritonic case, the diatomic molecule is embedded in a strongly coupled microcavity with variable host material with frequency $\omega_{M}$. The top panels, (a) and (b), correspond to HF, and the bottom panels, (c) and (d), show results for LiH. In (a) and (c), the collective light-matter interaction strength, $\Omega_{R}$, is fixed at 400 and 200 cm-1, respectively. The horizontal dashed gray and green lines indicate where the dissociation rates equal $k_{0}$ and $k_{c}$, respectively. The insets in (a) and (c) show zoomed-in views around the overtones $17\rightarrow 23$ for HF and $22\rightarrow 29$ for LiH. In (b) and (d), relative BIRD rates are shown for different Rabi frequencies ($\Omega_{R}$), depicting the dependency on the detuning between the host molecule and the diatomic transition energies.
Figure 5: Photon-weighted polariton DOS $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at $\omega_{M}=1401$ cm-1. The curves obtained with distinct $\Omega_{R}$ values illustrate its effect on the photon-weighted polariton DOS relevant for radiative processes mediated by polaritonic systems.
In Fig. 5, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=9.6$ $\mu$m and $\omega_{M}=v_{22\rightarrow 29}(\text{LiH})=1401$ cm-1. The case where $\Omega_{R}=0$ represents the weak coupling limit and is included here to directly compare $D_{P}(\omega)$ with $D_{C}(\omega)$. It can be seen from Fig. 5, the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap. Most remaining features, e.g., the oscillations in $D_{P}(\omega)$ with $\omega$, are also observed in weak coupling and can be understood from Eq. 8. The different curves in Fig. 5 show how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ grows with increasing $\Omega_{R}$. Therefore, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting. This becomes more likely for larger values of $\Omega_{R}$.
While the singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate, our results demonstrate that this singularity does not lead to unboundedly fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting. Therefore, our results obtained as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here. In experiments, imperfections in the electromagnetic device and the strongly coupled material smooth out the photon-weighted DOS near the stopgap. Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder Satanin et al. (2005). The introduction of a UV cutoff and loss pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig. 4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc.). At the same time, the reported theoretical upper bounds on diatomic BIRD rates (Fig. 4) remain valid. Finally, we emphasize that the same qualitative behavior is observed for polariton-mediated BIRD in LiH and HF. Since these chemical species have significantly different energetics, we expect the main trends reported here also apply to other diatomic and polyatomic molecules.
IV Conclusions
This work presented a detailed analysis of BIRD in infrared microcavities. Our analysis revealed mechanisms for controlling dissociation phenomena in microcavities, particularly by manipulating the electromagnetic density of states (DOS) in weak and strong light-matter coupling regimes. Overtones play a critical role in the dissociation process under both weak and strong coupling cases, as the corresponding transitions act as rate-limiting steps for the dissociative process. Therefore, the radiative dissociation rate changes induced by a microcavity are essentially tied to how the photonic density of states is modified around specific overtone transitions.
We reported microcavity-assisted dissociation rate enhancements of $\mathcal{O}(1)$ in the weak coupling regime. In microcavities where $L_{C}\leq 5~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates. As $L_{C}$ increases beyond a few microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones. Further increases in $L_{C}$ lead to BIRD rates approaching the free-space limit.
In the strong coupling regime, where polaritons mediate BIRD, we reported $\mathcal{O}(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process. This enhancement is due to the emergence of a singularity in the photon-weighted polariton DOS, which massively increases transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below. The singularity is attributed to the lack of a high-energy cutoff for the electromagnetic field and the lossless character of the strongly coupled system. Nevertheless, while the singular behavior of the photon-weighted polariton DOS is absent from experiments, we have argued its effects are expected to persist to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder. Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced changes in diatomic thermal radiative dissociation.
Our work focused on the dissociation of diatomic molecules for simplicity. The main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly and are expected to be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.
Acknowledgments. R.F.R. acknowledges support from the ACS Petroleum Research Fund under project PRF-66399-DNI6, NSF CAREER award Grant No. CHE-2340746, and startup funds provided by Emory University.
Author Contributions. G. J. R. A. and E. S. contributed equally to this work.

### v2

@@
 Vibrational strong light-matter coupling offers a promising approach for controlling chemical reactivity with infrared microcavities.
-This study explores the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes.
+While recent research has examined potential mechanisms for this phenomenon, many important questions remain, including what type of reactions can be modified and to what extent this modification can be achieved.
+In this study, we explore the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes.
 Using a Master equation approach, we simulate the effects of infrared field confinement and vibrational strong coupling on BIRD rates for diatomic molecules.
 We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process.
-Our findings reveal conditions for significant enhancement and mild suppression of radiative dissociation, establishing upper bounds for BIRD rates under weak and strong coupling.
-These results provide new strategies and limitations for controlling reactive processes with infrared resonators.
+Our findings outline the conditions under which significant enhancement or mild suppression of BIRD rates can be achieved, offering insights into the practical limitations and new strategies for controlling chemistry within infrared resonators.
 I   Introduction
@@
 Conventional methods for manipulating reaction rates involve varying temperature, pressure, or catalyst composition.
-An unconventional approach for controlling thermal reaction kinetics has been recently reported Thomas et al.
-(2016, 2019); Hirai et al.
-(2020); Ahn et al.
-(2023) employing infrared (IR) microcavities Kavokin et al.
-(2017) consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations Nagarajan, Thomas, and Ebbesen (2021).
-When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device Vinogradov (1992); Shalabney et al.
-(2015); Simpkins et al.
-(2015).
-Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction ratesGalego et al.
-(2019); Campos-Gonzalez-Angulo, Ribeiro, and Yuen-Zhou (2019); Zhdanov (2020); Wang, Flick, and Yelin (2022); Lindoy, Mandal, and Reichman (2022); Schäfer et al.
-(2022); Sun and Vendrell (2022); Ahn et al.
-(2023); Montillo Vega, Ying, and Huo (2024); Moiseyev (2024).
-Still, several fundamental questions remain Simpkins, Dunkelberger, and Owrutsky (2021); Campos-Gonzalez-Angulo et al.
-(2023), including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted.
-In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation Dunbar (2004) under weak and strong coupling with a suitable material.
-Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation Perrin (1970).
-However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism Lindemann et al.
-(1922); Daniels (1928); King and Laidler (1984).
-Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressuresThoelmann, Tonner, and McMahon (1994); Dunbar et al.
-(1995); Schnier et al.
-(1996).
-Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers Dunbar (2004).
-BIRD rates are heavily influenced by the thermal radiation spectrum.
-Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different Kavokin et al.
-(2017), microcavity BIRD is expected to proceed with different rates compared to free space BIRD.
-However, this difference’s order of magnitude, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime.
+An unconventional approach for controlling thermal reaction kinetics has been recently reported [1, 2, 3, 4] employing infrared (IR) microcavities [5] consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations [6].
+When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device [7, 8, 9].
+Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction rates[10, 11, 12, 13, 14, 15, 16, 4, 17, 18, 19].
+Still, several fundamental questions remain [20, 21], including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted.
+In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation [22] under weak and strong coupling with a suitable material.
+Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation [23].
+However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism [24, 25, 26].
+Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressures[27, 28, 29].
+Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers [22].
+The thermal radiation spectrum heavily influences BIRD rates.
+Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different [5], microcavity BIRD is expected to proceed with different rates compared to free space BIRD.
+However, the extent of this change, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime remain open questions.
 Our study provides answers to these questions.
-Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but our qualitative conclusions are generalizable to polyatomic systems.
+Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but several qualitative conclusions are generalizable to polyatomic systems.
 The primary aim of this study is to characterize thermal radiative dissociation in microcavities.
-We employ a Master equation treatment to investigate blackbody infrared radiative dissociation for diatomic molecules in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material.
-Our results provide upper bounds for enhancement and suppression of BIRD in weakly and strongly coupled microcavities and identify overtone transitions as key ingredients for the observed modulation of BIRD.
+We employ a Master equation treatment to investigate this process in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material.
+Our results give upper bounds for BIRD enhancement and suppression in weakly and strongly coupled microcavities and identify overtone transitions as crucial ingredients for the observed modulation of BIRD.
 In Sec.
 II, we describe our methodology, Sec.
-III provides our main results and related discussion, and Sec.
+III provides the main results and related discussion, and Sec.
 IV summarizes our main results and conclusions.
@@
 (d) and (e): A scheme illustrating an HF molecule inside a planar cavity with the host material strongly coupled to the cavity and the dispersion of polariton states, respectively.
-II.1 Master Equation for diatomic BIRD
-Our analysis of BIRD in microcavities employs the same tools successfully applied to investigate this process in free space Dunbar (2004); Dunbar et al.
-(1995); Stevens et al.
-(2002); Dunbar and McMahon (1998).
-These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold.
-Rotational degrees of freedom are ignored since at the temperatures of interest to BIRD, the rotational dynamics is essentially classical, only introducing a nonzero linewidth to the vibrational absorption and emission spectrum.
-We discuss further implications of this condition in the context of our results below.
-The rate of change of the population of the $i$th vibrational state ($N_{i}$) is given by
+Master Equation for diatomic BIRD
+Our analysis of BIRD in microcavities employs similar tools to those successfully applied to investigate this process in free space [28, 30, 31, 22].
+These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold defined below.
+Rotational degrees of freedom are ignored since they are weakly coupled to vibrational transitions, and the rotational dynamics are essentially classical at the temperatures of interest to BIRD.
+We also neglect Doppler and lifetime broadening throughout.
+The impact of nonzero vibrational linewidths on the presented results is described in Results and Discussion.
+For the sake of simplicity, we assumed the thermalization of the background radiation is faster than all considered radiative transitions.
+The rate of change of the population in the $i$th vibrational state ($N_{i}$) is given by
 $\frac{dN_{i}(t)}{dt}=\sum_{j\neq i}^{i_{\text{max}}}\left[k_{ij}N_{j}(t)-k_{ji}N_{i}(t)\right]-k_{\text{loss}}N_{i}(t)\delta_{i,i_{\text{max}}}\;,$ (1)
-where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ refers to the highest energy bound state (Fig.
+where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ is the highest energy bound state (Fig.
 1a).
-We assume that the state $i_{\text{max}}$ undergoes irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$.
-For the sake of simplicity, we have assumed the thermalization of the background radiation is faster than all considered radiative transitions.
+This state is only loosely bound and is assumed to undergo irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$.
 Equation 1 can be written in matrix notation as
@@
 where N is a column vector containing the vibrational populations and $\mathbf{J}$ is the transition matrix.
-The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction Valance and Schlag (1966); Gilbert and Smith (1990).
-II.2 Radiative transition rates
-Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule Fermi (1932); Sakurai and Napolitano (2017)) and written in terms of Einstein coefficients in free space Einstein (1917); Cohen-Tannoudji, Diu, and Laloë (2019) and their generalizations in a weakly coupled and polaritonic microcavity.
+The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction [32, 33] of the described mechanism.
+Note that this mechanism does not consider dissociation from intermediate vibrational levels ($\nu<i_{\text{max}}$).
+Thus, our results are expected to be valid for BIRD promoted by multiphoton dissociation through a particular doorway state, which is the highest-energy bound state in our case.
+Radiative transition rates
+Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule [34, 35]) and written in terms of Einstein coefficients in free space [36, 37] and their generalizations in a weakly coupled and polaritonic microcavity.
 Rotational dynamics is much faster than thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
-In free space, the spontaneous emission rate is given by the Einstein $A$ coefficientEinstein (1917); Cohen-Tannoudji, Diu, and Laloë (2019).
+In free space, the spontaneous emission rate is given by the Einstein $A$ coefficient[36, 37].
 In the case where spontaneous emission induces the $i\rightarrow j$ transition with $i>j$ this coefficient is given by
@@
 $\rho_{0}(\omega,T)=\hbar\omega~{}n_{\text{BE}}(\omega,T)D_{0}(\omega),$ (5)
-where $n_{\text{BE}}(\omega,T)$ is the Bose-Einstein distribution thermal occupation number $n_{\text{BE}}(\omega,T)=1/[\text{exp}(\hbar\omega/k_{B}T)-1]$Bose (1924) and $k_{B}$ is the Boltzmann constant.
+and $n_{\text{BE}}(\omega,T)$ is the Bose-Einstein distribution thermal occupation number $n_{\text{BE}}(\omega,T)=1/[\text{exp}(\hbar\omega/k_{B}T)-1]$[38] and $k_{B}$ is the Boltzmann constant.
 The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$.
 We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic molecules is given by the Morse potential [Fig.
-1(a)]Morse (1929).
+1(a)][39].
 The bound wave functions associated with the Morse potential are also employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
-$\mu_{ij}$ $=\int\psi_{i}(r)\mu(r)\psi_{j}(r)\mathrm{d}r,$ (6)
-where $\psi_{i}(r)$ represents the $i$th vibrational level wave function, and the dipole function $\mu(r)$ models the nonlinear change in the (electronic ground-state) molecular electrical dipole with the bond length $r$ Bishop (1990).
+$\mu_{ij}$ $=\int\psi^{*}_{i}(r)\mu(r)\psi_{j}(r)\mathrm{d}r,$ (6)
+where $\psi_{i}(r)$ represents the $i$th vibrational level wave function, and the dipole function $\mu(r)$ models the nonlinear change in the (electronic ground-state) molecular electrical dipole with the bond length $r$ [40].
 We investigated the IR radiative dissociation of hydrogen fluoride, HF, and lithium hydride, LiH.
-The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-1 were obtained from previous work on laser-induced HF dissociationKaluža et al.
-(1994).
-The LiH dipole function was interpolated from ab initio quantum chemical computations (see Supplementary Information).
+The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-4 were obtained from previous work on laser-induced HF dissociation[41].
+The LiH dipole function was interpolated from ab initio quantum chemical computations (CCSD(T)/aug-cc-pVTZ, see Supplementary Information for further details).
 Morse potential parameters include the equilibrium bond length ($r_{e}$), electronic ground-state dissociation energy ($D_{e}$), and harmonic frequency ($\omega_{e}$) of each molecule and are listed in Table 1.
-Table 1: Parameters used for Morse potential of HF Stine and Noid (1979); Guldberg and Billing (1991); Gross, Neuhauser, and Rabitz (1993); Kaluža et al. (1994) and LiH Irikura (2007).
+Table 1: Parameters used for Morse potential of HF [42, 43, 44, 41] and LiH [45].
 Constant HF LiH
@@
 $m_{r}$ (a.u) 0.95 0.88
-Note that the examined diatomic species have severely reduced room temperature BIRD rates relative to typical experiments involving polyatomic ions Dunbar and McMahon (1998); Dunbar et al.
-(1995) due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation.
+The examined dissociation processes (via the most energetic bound state) have reduced room temperature rates relative to typical BIRD experiments involving polyatomic ions [27, 46, 29, 22] due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation proceeding via the highest energy level.
 This feature motivated our choice of examining diatomic BIRD at temperatures $T=\mathcal{O}(\hbar\omega_{e}/k_{\text{B}})$, an order of magnitude greater than room temperature.
-Such temperatures lead to numerically stable BIRD rates in the various regimes considered in this work.
-Despite their largely different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior.
-This independence to microscopic details demonstrates the robustness of our analysis.
+The SI shows that the trends discussed below hold for a wide interval of temperatures.
+Despite their different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior.
 The transition rates $k_{ij}$ are proportional to the electromagnetic density of states $D(\omega)$, suggesting BIRD rates can be controlled by placing the reactive molecule inside a microcavity.
@@
 1(b)], the frequencies of the electromagnetic modes are given by
-$\omega_{\text{C}}(m,\textbf{q})=c\sqrt{q^{2}+\frac{m^{2}\pi^{2}}{L_{C}^{2}}}\;,$ (7)
+$\omega_{\text{C}}(m,q)=c\sqrt{q^{2}+\frac{m^{2}\pi^{2}}{L_{C}^{2}}}\;,$ (7)
 where $\mathbf{q}=(q_{x},q_{y})\in\mathbb{R}^{2}$ is the in-plane projection of the mode wave vector, $q^{2}=q_{x}^{2}+q_{y}^{2}$, and $m$ is an integer quantum number that identifies the longitudinal wave vector $k_{z}=m\pi/L_{C}$ and thus the photonic energy bands in Fig.
 1(c).
-For transverse-magnetic (TM) modes $m=0,1,2,...$ whereas $m=1,2,3...$ for transverse-electric (TE) modes Zoubi and La Rocca (2005); Jackson (2021).
+For transverse-magnetic (TM) modes $m=0,1,2,...$ whereas $m=1,2,3...$ for transverse-electric (TE) modes [47, 48].
 Including the TM0 mode here is crucial, as otherwise, transitions with energy below the cavity cutoff would be impossible.
-Using Eq.
-7, it can be shown (see Supporting Information) that the spatially-averaged electromagnetic DOS inside a microcavity in the weak coupling regime can be expressed by Barnes, Horsley, and Vos (2020)
+Using Eq. 7, it can be shown (see Supporting Information) that the spatially-averaged electromagnetic DOS inside a microcavity in the weak coupling regime can be expressed by [49]
 $D_{C}(\omega)=\frac{\omega}{\pi c^{2}L_{C}}\left\lfloor\frac{\omega L_{C}}{\pi c}\right\rfloor+\frac{\omega}{2\pi c^{2}L_{C}}\;,$ (8)
@@
 1(d) and (e).
-As we show in the SI, the formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
-The Hamiltonian of the strongly coupled molecular subsystem and the microcavity is given in the Power-Zienau-Woolley (dipole) gauge Power and Zienau (1959); Woolley (1971)
+The formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD (see SI) by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
+The Hamiltonian of the strongly coupled molecular subsystem and the microcavity is given in the Power-Zienau-Woolley (dipole) gauge [50, 51]
 $H=H_{\text{SM}}+H_{\text{L}}+H_{\text{INT}}\;,$ (9)
@@
 […]
-$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-m\pi c/L_{C}],$ (11)
-where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$ (see SI)
+$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-\frac{m\pi c}{L_{C}}],$ (11)
+where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$
 $P_{C}(\omega)=\frac{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}}{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}+\omega^{2}_{C}(\omega)\Omega_{R}^{2}},$ (12)
@@
 In this limit, both the photon content $P_{C}(\omega)$ and the group velocity $v_{g}(\omega,q)$ approach zero, but the ratio $P_{C}(\omega)/v_{g}(\omega,q)\rightarrow\infty$ resulting in $D_{P}(\omega)$ diverging to positive infinity.
-This singularity is due to the formation of the well-known stopgap regionLitinskaya and Agranovich (2009); Huang et al.
-(2003); Gómez Rivas et al.
-(2004); Todorov and Sirtori (2012) where $D_{P}(\omega)=0$.
-As previous work shows Satanin et al.
-(2005), introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$.
+This singularity is due to the formation of the well-known stopgap region[59, 60, 61, 62] where $D_{P}(\omega)=0$.
+As previous work shows [63], introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$.
 Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on the type of disorder and coherence loss mechanism.
-This motivates our choice to work with the formally exact expression given in Eq.
-11 and interpret our results as providing upper bounds for the effect of polaritons on BIRD.
+This motivates our decision to use the formally exact expression presented in Eq.
+11 and interpret our results as upper bounds on the effect of polaritons on BIRD.
 III   Results and Discussion
-III.1 Microcavity-assisted blackbody infrared radiative dissociation
+Microcavity-assisted blackbody infrared radiative dissociation
 Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity.
-The tunable electromagnetic environment afforded by the microcavity allows suggests BIRD rates can be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the change induced by the microcavity in the photon DOS.
+The tunable electromagnetic environment afforded by the microcavity allows BIRD rates to be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the modified photon DOS.
 In Fig.
 2, BIRD rates in the weak coupling regime are shown relative to the free space rate ($k_{0}=2.70\times 10^{-6}$ s-1 for HF at $T/\omega_{e}=0.67$ corresponding to $T=4000~{}\text{K}$ and $k_{0}=1.64\times 10^{-5}$ s-1 for LiH at $T/\omega_{e}=0.99$ corresponding to $T=2000~{}\text{K}$).
-For small microcavity lengths ($L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH), a consistent mild enhancement of the BIRD rate is associated with an overall increase in the microcavity electromagnetic DOS relative to free space over a wide frequency range covering most fundamental transitions and relevant overtones of HF and LiH.
+For small microcavity lengths, i.e., $L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH, a mild but consistent enhancement of BIRD rates is observed.
+At these lengths, an overall increase in the microcavity electromagnetic DOS over a wide frequency range covering several important transitions is associated with the observed enhancement.
 The ratios of microcavity BIRD rates to free space BIRD rates quickly approach 1.0 as $L_{C}$ increases beyond a few microns.
@@
 The complex oscillatory pattern in the BIRD rates relative to free space shown in Fig.
-2 arises from the interplay of transitions enhanced or suppressed depending on $L_{C}$.
+2 arises from the interplay of discrete vibrational transitions enhanced or suppressed depending on $L_{C}$.
 As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
@@
 3.
-To interpret the origin of the oscillatory behavior in Fig.
-2 and identify the most consequential radiative transitions involved in BIRD of HF and LiH, we conducted a sensitivity analysis of the dissociation rate.
-This measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details).
+We conducted a sensitivity analysis of the dissociation rate to interpret the origin of the oscillatory behavior in Fig.
+2 and identify the most consequential radiative transitions involved in the BIRD of HF and LiH.
+This analysis measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details).
 For HF dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 23$, with $i\in\{16,17,19\}$ (in the relevant 1500 -- 5000 cm-1 frequency range) are the most consequential for the BIRD rate.
 Similarly, for LiH, the relevant transitions are $i\rightarrow 29$, with $i\in\{21,22,23,25\}$.
-In both cases, the identified overtones take the vibrational mode to the highest energy bound state (23 for HF and 29 for LiH) from which dissociation rapidly occurs.
+In both cases, the identified overtones take the vibrational mode to the highest energy-bound state (23 for HF and 29 for LiH), from which dissociation rapidly occurs.
 We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths.
-Figure 3: Microcavity photon density of states normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig.
+See Supporting Information for a brief discussion of results obtained without overtone transitions.
+Figure 3: Microcavity photon DOS normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig.
 2.
-The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color-coding to distinguish between suppression (red) and enhancement (green).
-Panel (a) shows the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced microcavity photon DOS relative to free space at the highlighted overtones.
+The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color coding to distinguish between suppression (red) and enhancement (green).
+Panel (a) shows that the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced relative DOS at the highlighted overtones.
 Panel (b) shows the mild enhancement of HF BIRD in a microcavity with $L_{C}=3.21~{}\mu\text{m}$ can be ascribed to the greater microcavity photon DOS at the frequencies of the highlighted overtones.
 Panels (c) and (d) show the analogous scenario for LiH, where BIRD is observed to be suppressed at $L_{C}=9.4~{}\mu\text{m}$ but enhanced at $L_{C}=9.6~{}\mu\text{m}$, primarily due to the variation of the microcavity photon density of states at the selected $i\rightarrow 29$ overtone frequencies.
-Equipped with the knowledge of the most relevant transitions for BIRD of HF and LiH, we can readily explain the microscopic origin of the intricate pattern in Fig.
+Equipped with the knowledge of the most relevant transitions for the investigated BIRD of HF and LiH via the most energetic bound state, we can readily explain the microscopic origin of the intricate pattern in Fig.
 2 by considering how the microcavity DOS changes around the previously mentioned overtones.
@@
 In Fig.
-3a ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig.
+3(a) ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig.
 2.
-In contrast, when the microcavity length is slightly increased to $L_{C}=3.21$ $\mu$m (Fig.
-3(b) the photon DOS is enhanced at $\omega$ corresponding to the $16\rightarrow 23$ and $19\rightarrow 23$ transitions, and we observe a corresponding leap in the BIRD rate in Fig.
+In contrast, when the microcavity length slightly increases to $L_{C}=3.21$ $\mu$m [Fig.
+3(b)] the photon DOS is enhanced at the $16\rightarrow 23$ and $19\rightarrow 23$ transitions frequencies, and we observe a corresponding leap in the BIRD rate in Fig.
 2.
@@
 3(d) shows that at $L_{C}=9.6~{}\mu\text{m}$ most of the relevant LiH overtones are enhanced, thus explaining the sudden increase in BIRD rate at $L_{C}=9.6~{}\mu\text{m}$ relative to $L_{C}=9.4~{}\mu\text{m}$.
-As we conclude this subsection, we note that suppression effects are nearly negligible in Fig.
+Introducing a finite linewidth to the considered radiative transitions would smooth out the oscillations observed in Fig.
+2 and dampen their amplitude.
+Consequently, the weakly coupled microcavity effect on BIRD would be significant only at lengths less than a few microns.
+Note that suppression effects are generally minor in Fig.
 2 because the dissociative process has many pathways.
-Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always nearly negligible as the molecular system explores alternative dissociation pathways.
-Therefore, our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
-Strictly speaking, we have shown that these conditions rigorously apply to diatomic BIRD, but they are likely to be even more relevant in polyatomic processes.
-III.2 Polariton-assisted blackbody infrared radiative dissociation
+Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always negligible as the molecular system explores alternative dissociation pathways.
+Our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
+Polariton-assisted blackbody infrared radiative dissociation
 To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes.
-Figure 1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photonfrequency-momentum dispersion relation [Fig.
+Fig.
+1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photon frequency-momentum dispersion relation [Fig.
 1 (e)] and the electromagnetic DOS.
-We explore the implications for diatomic BIRD below.
-Figure 4 presents the ratio of polariton-assisted HF and LiH BIRD rates to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$.
-These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 at microcavity length $L_{C}$ fixed at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH, respectively.
-As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime.
-We observe prominent peaks for both HF and LiH where the relative dissociation rate is notably enhanced.
-Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$, where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH.
-Conversely, as seen in the inset of Figs.
-4(a) and (c), when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed.
-This happens because when $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the relevant overtone transitions fall into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig.
-1e).
-Nonetheless, the suppression is small since, as discussed above, BIRD may proceed via other pathways.
-The effect of a variation of collective coupling strengths ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig.
-4(b) and (d).
-When $\omega_{M}-\omega_{i\rightarrow j}$ is around 1-10 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates.
-However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively.
-This plateau appears to be insensitive to the magnitude of the Rabi splitting.
-It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig.
-5.
+We explore the implications for diatomic BIRD via the most energetic bound state below.
 Figure 4: Ratio of polariton-assisted BIRD rates to free space BIRD rates.
@@
 In (b) and (d), relative BIRD rates are shown for different Rabi frequencies ($\Omega_{R}$), depicting the dependency on the detuning between the host molecule and the diatomic transition energies.
-Figure 5: Photon-weighted polariton DOS $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at $\omega_{M}=1401$ cm-1.
-The curves obtained with distinct $\Omega_{R}$ values illustrate its effect on the photon-weighted polariton DOS relevant for radiative processes mediated by polaritonic systems.
+Fig.
+4 presents the ratio of polariton-assisted BIRD to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$.
+These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 by fixing $L_{C}$ at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH respectively.
+As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime.
+We observe prominent peaks for HF and LiH where the relative dissociation rate is notably enhanced.
+Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$.
+Here, we focus on the case where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH.
+In contrast, the inset of Figs.
+4(a) and (c) show that when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed.
+This suppression occurs because as $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the overtone transition falls into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig.
+1e).
+However, the BIRD suppression is relatively small since, as discussed earlier, the reaction may still proceed through alternative pathways.
+The effect of the collective coupling strength ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig.
+4(b) and (d).
+When $\omega_{M}-\omega_{i\rightarrow j}$ is around 0.01-1 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates.
+However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively.
+This plateau appears to be insensitive to the magnitude of the Rabi splitting.
+It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig.
+5.
+Figure 5: Photon-weighted polariton density of states $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with length $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at frequency $\omega_{M}=1401$ cm-1.
+The curves show the effect of different Rabi splitting values ($\Omega_{R}$) on the photon-weighted polariton DOS, which is relevant for radiative processes mediated by polaritonic systems.
+As $\Omega_{R}$ increases, the frequency range below $\omega_{M}$, where the photon-weighted polariton DOS significantly exceeds that of free space, widens.
+This trend explains the impact of the collective light-matter interaction strength on the observed polariton-assisted bond infrared dissociation (BIRD) enhancement shown in Fig.
+4.
 In Fig.
@@
 It can be seen from Fig.
-5, the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap.
+5 that the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap.
 Most remaining features, e.g., the oscillations in $D_{P}(\omega)$ with $\omega$, are also observed in weak coupling and can be understood from Eq.
 8.
-The different curves in Fig.
-5 show how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ grows with increasing $\Omega_{R}$.
-Therefore, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting.
-This becomes more likely for larger values of $\Omega_{R}$.
-While the singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate, our results demonstrate that this singularity does not lead to unboundedly fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting.
-Therefore, our results obtained as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here.
+The various curves in Fig.
+5 illustrate how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ expand as $\Omega_{R}$ increases.
+As a result, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting.
+The singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate.
+However, our results demonstrate that this singularity does not lead to arbitrarily fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting.
+Therefore, the BIRD rates reported as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here.
 In experiments, imperfections in the electromagnetic device and the strongly coupled material smooth out the photon-weighted DOS near the stopgap.
-Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder Satanin et al.
-(2005).
-The introduction of a UV cutoff and loss pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig.
-4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc.).
-At the same time, the reported theoretical upper bounds on diatomic BIRD rates (Fig.
-4) remain valid.
-Finally, we emphasize that the same qualitative behavior is observed for polariton-mediated BIRD in LiH and HF.
-Since these chemical species have significantly different energetics, we expect the main trends reported here also apply to other diatomic and polyatomic molecules.
+Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder [63].
+The introduction of a UV cutoff, loss, and dephasing pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig.
+4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc).
+Including a finite linewidth for the diatomic vibrational transitions relaxes the strict resonance requirement discussed above ($\omega_{M}\rightarrow\omega_{i\rightarrow j}$), and the overall enhancement will depend on the spectral overlap between $\omega_{i\rightarrow j}$ and $D_{P}(\omega)$.
+Since the molecular linewidth and cavity imperfections are expected to be limiting factors, it is clear that our reported results represent theoretical upper bonds for polariton-assisted BIRD rates for the reaction mechanism investigated here.
+Before ending, we emphasize our observation of the same qualitative behavior for polariton-mediated BIRD in LiH and HF.
+Since these chemical species have significantly different energetics, we anticipate the main trends reported in this work also apply to other diatomic and polyatomic molecules, particularly in cases where dissociation occurs through a specific doorway state.
 IV   Conclusions
@@
 Our analysis revealed mechanisms for controlling dissociation phenomena in microcavities, particularly by manipulating the electromagnetic density of states (DOS) in weak and strong light-matter coupling regimes.
-Overtones play a critical role in the dissociation process under both weak and strong coupling cases, as the corresponding transitions act as rate-limiting steps for the dissociative process.
-Therefore, the radiative dissociation rate changes induced by a microcavity are essentially tied to how the photonic density of states is modified around specific overtone transitions.
-We reported microcavity-assisted dissociation rate enhancements of $\mathcal{O}(1)$ in the weak coupling regime.
+Overtones play a critical role in the dissociation process under both weak and strong coupling cases, as the corresponding transitions act as rate-limiting steps for the dissociative process occurring via the highest energy-bound state.
+The radiative dissociation rate changes induced by a microcavity are essentially tied to how the photonic density of states is modified around specific overtone transitions driving the system into the dissociating (doorway) state.
+We reported microcavity-assisted dissociation rate enhancements of $O(1)$ in the weak coupling regime.
 In microcavities where $L_{C}\leq 5~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates.
 As $L_{C}$ increases beyond a few microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones.
-Further increases in $L_{C}$ lead to BIRD rates approaching the free-space limit.
-In the strong coupling regime, where polaritons mediate BIRD, we reported $\mathcal{O}(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process.
+Further increases in $L_{C}$ causes the BIRD rates to approach the free-space limit.
+In the strong coupling regime, where polaritons mediate BIRD, we reported $O(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process.
 This enhancement is due to the emergence of a singularity in the photon-weighted polariton DOS, which massively increases transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below.
@@
 Nevertheless, while the singular behavior of the photon-weighted polariton DOS is absent from experiments, we have argued its effects are expected to persist to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder.
-Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced changes in diatomic thermal radiative dissociation.
-Our work focused on the dissociation of diatomic molecules for simplicity.
-The main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly and are expected to be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.
-Acknowledgments. R.F.R. acknowledges support from the ACS Petroleum Research Fund under project PRF-66399-DNI6, NSF CAREER award Grant No. CHE-2340746, and startup funds provided by Emory University.
+Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced diatomic thermal radiative dissociation changes via the most energetic bound state.
+Our work focused on how optical microcavities can modify BIRD reactions of diatomic molecules via a particular mechanism.
+Despite focusing on these specific systems, the main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly.
+We expect the analysis introduced here will be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.
+Acknowledgments.
+This work was supported by the donors of ACS Petroleum Research Fund under Doctoral New Investigator Grant 66399-DNI6.
+R.F.R.
+served as Principal Investigator on ACS PRF-66399-DNI6 that provided support for E.
+S.
+R.F.R also acknowledges support from NSF CAREER award Grant No.
+CHE-2340746, and startup funds provided by Emory University.
 Author Contributions. G. J. R. A. and E. S. contributed equally to this work.

### v3

@@
 In this study, we explore the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes.
-Using a Master equation approach, we simulate the effects of infrared field confinement and vibrational strong coupling on BIRD rates for diatomic molecules.
+Using a Master equation approach, we simulate the effects of infrared field confinement and polariton formation on BIRD rates for diatomic molecules weakly coupled to the radiation field.
 We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process.
-Our findings outline the conditions under which significant enhancement or mild suppression of BIRD rates can be achieved, offering insights into the practical limitations and new strategies for controlling chemistry within infrared resonators.
+Our findings outline conditions under which significant enhancement or mild suppression of BIRD rates can be achieved, offering insights into practical limitations and new strategies for controlling chemistry within infrared resonators.
 I   Introduction
@@
 When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device [7, 8, 9].
-Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction rates[10, 11, 12, 13, 14, 15, 16, 4, 17, 18, 19].
+Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction rates[10, 11, 12, 13, 14, 15, 16, 4, 17, 18, 19], and experimental progress towards the characterization of gas-phase reactivity under strong light-matter coupling has also been reported [20, 21].
 Still, several fundamental questions remain[22, 23], including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted.
@@
 Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressures[29, 30, 31].
-Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers [22].
+Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several systems, including ion clusters, transition-metal complexes, and biopolymers [24].
 The thermal radiation spectrum heavily influences BIRD rates.
@@
 However, the extent of this change, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime remain open questions.
-Our study provides answers to these questions.
-Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but several qualitative conclusions are generalizable to polyatomic systems.
-The primary aim of this study is to characterize thermal radiative dissociation in microcavities.
-We employ a Master equation treatment to investigate this process in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material.
-Our results give upper bounds for BIRD enhancement and suppression in weakly and strongly coupled microcavities and identify overtone transitions as crucial ingredients for the observed modulation of BIRD.
+In this study, we address these questions by examining the dissociation of diatomic molecules driven by multiphoton or multipolariton absorption in microcavities.
+The qualitative conclusions drawn from this analysis are expected to be generalizable to polyatomic systems.
+The primary aim of this work is to characterize thermal radiative dissociation in microcavities.
+Whereas prior studies employed hierarchical equations of motion and analytical rate theories based on Fermi’s golden rule to explore energy-diffusion-limited reactivity modeled by double-well potentials [14, 16, 32, 33, 34, 17], here we employ a Pauli master equation to investigate irreversible bond breaking with a Morse potential in a scenario where the microcavity or polaritons act solely as passive modifiers of the electromagnetic environment.
+Our methodology further differs by its microscopically detailed treatment of mechanical and electrical anharmonicity inherent to reactivity processes, and by its inclusion of the coupling of both fundamental and overtone vibrational transitions to the modified photon density of states, thereby providing a detailed description of the microcavity effect on the multiple reactive pathways available in BIRD.
+We employ a Pauli Master equation to investigate BIRD in (i) free-space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material.
+Our results give upper bounds for BIRD enhancement and suppression in weakly and strongly coupled microcavities with perfect mirrors (see SI Sec.
+8 for a discussion of leaky microcavities) and identify overtone transitions as crucial ingredients for the observed modulation of BIRD.
 In Sec.
@@
 II   Methods
-Figure 1: (a) Illustration of the Morse potential describing the vibrational states and transitions of the HF molecule.
-(b) and (c): A scheme illustrating an HF molecule inside a planar cavity and its dispersion under the weak coupling regime, respectively.
-(d) and (e): A scheme illustrating an HF molecule inside a planar cavity with the host material strongly coupled to the cavity and the dispersion of polariton states, respectively.
-Master Equation for diatomic BIRD
-Our analysis of BIRD in microcavities employs similar tools to those successfully applied to investigate this process in free space [28, 30, 31, 22].
-These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold defined below.
+Figure 1: (a) Illustration of the Morse potential describing the vibrational states and considered transitions of the NaLi molecule.
+(b) Illustration of a diatomic molecule inside a high quality planar microcavity with longitudinal length $L_{C}$.
+(c) Dispersion curves of transverse-electric (TE) and transverse-magnetic (TM) microcavity modes.
+(d) Scheme illustrating a reactive diatomic molecule inside a planar resonator strongly coupled to a host material system with significant oscillator strength and transition frequency $\omega_{M}$ (dashed line at $\omega=0.05~{}\text{eV}$).
+In this setup, the diatomic dissociation is mediated by absorption of thermal polaritons.
+(e) Illustration of polariton dispersion curves.
+II.1      Pauli Master Equation for diatomic BIRD
+Our analysis of BIRD in microcavities employs the same kinetic framework that has proven successful for BIRD studies in free space [30, 35, 36, 24].
+Specifically, we propagate the populations of bound vibrational levels with a Pauli master equation[37] including transition rates obtained from Fermi’s golden rule [38].
+This choice is justified because the infrared absorption and emission events of interest ($>$ 100 ns) are orders of magnitude slower than rovibrational decoherence, which randomizes vibrational coherences within on a much faster timescale.
+The resulting separation of time scales renders the diatomic reduced density matrix effectively diagonal (in the Morse potential Hamiltonian eigenstate basis) well before any population transfer occurs, allowing the Liouville–von Neumann equation to be coarse-grained to the Pauli form [37, 39, 40].
+We neglect collisional activation and assume that dissociation is driven exclusively by photon absorption, consistent with the conditions of BIRD experiments[30, 24].
+In such experiments, the use of dilute gas samples at low pressures ensures that intermolecular collisions are negligible.
+External fields are applied to trap and stabilize molecular trajectories within the interior of the reaction vessel.
+This prevents molecules from exchanging energy with walls and allows radiative processes to dominate [41, 42, 43].
 Rotational degrees of freedom are ignored since they are weakly coupled to vibrational transitions, and the rotational dynamics are essentially classical at the temperatures of interest to BIRD.
 We also neglect Doppler and lifetime broadening throughout.
-The impact of nonzero vibrational linewidths on the presented results is described in Results and Discussion.
-For the sake of simplicity, we assumed the thermalization of the background radiation is faster than all considered radiative transitions.
-The rate of change of the population in the $i$th vibrational state ($N_{i}$) is given by
+For the sake of simplicity, we assumed that the thermalization of the background radiation is faster than all the considered radiative transitions.
+We analyzed microcavity effects on BIRD occurring via two distinct mechanisms.
+In the first, denoted, dissociation via the doorway state, the dissociation proceeds only via the highest-energy bound state (the doorway state).
+In this case, the rate of change of the population in the $i$th vibrational state ($N_{i}$) is given by
 $\frac{dN_{i}(t)}{dt}=\sum_{j\neq i}^{i_{\text{max}}}\left[k_{ij}N_{j}(t)-k_{ji}N_{i}(t)\right]-k_{\text{loss}}N_{i}(t)\delta_{i,i_{\text{max}}}\;,$ (1)
@@
 1a).
-This state is only loosely bound and is assumed to undergo irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$.
+As mentioned, this state is only loosely bound and is assumed to undergo irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$.
 Equation 1 can be written in matrix notation as
@@
 The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction [44, 45] of the described mechanism.
-Note that this mechanism does not consider dissociation from intermediate vibrational levels ($\nu<i_{\text{max}}$).
-Thus, our results are expected to be valid for BIRD promoted by multiphoton dissociation through a particular doorway state, which is the highest-energy bound state in our case.
-Radiative transition rates
+Note that this mechanism does not consider dissociation from intermediate vibrational levels ($i<i_{\text{max}}$).
+Thus, our results are expected to be valid for BIRD promoted by multiphoton dissociation through a particular doorway state, which in our case is the highest-energy bound state.
+The second dissociation mechanism considered allows for direct radiative dissociation from any bound state of the diatomic molecule.
+Compared to Eq.
+1, this introduces a dissociation channel for each bound state, characterized by a loss rate $k^{i}_{\text{loss}}$.
+The corresponding population rate equations are given by
+$\frac{dN_{i}(t)}{dt}=\sum_{j\neq i}^{\nu_{\text{max}}}\left[k_{ij}N_{j}(t)-k_{ji}N_{i}(t)\right]-k^{i}_{\text{loss}}N_{i}(t).$ (3)
+The transition matrix $\mathbf{J}$ now includes contributions from bound-continuum decay rates in all diagonal elements.
+The presence of $k_{i}^{\text{loss}}$ for all $i$ ensures that the vibrational population can be irreversibly removed from each level due to dissociation.
+The matrix $\mathbf{J}$ remains positive-definite, and its smallest eigenvalue $\lambda_{0}$ continues to govern the long-time dynamics of the system.
+II.2      Radiative transition rates
 Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule [46, 47]) and written in terms of Einstein coefficients in free space [48, 49] and their generalizations in a weakly coupled and polaritonic microcavity.
-Rotational dynamics is much faster than thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
+Rotational dynamics is much faster than the thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
 In free space, the spontaneous emission rate is given by the Einstein $A$ coefficient[48, 49].
@@
 The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$.
-We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic molecules is given by the Morse potential [Fig.
+We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic is given by the Morse potential [Fig.
 1(a)][51].
-The bound wave functions associated with the Morse potential are also employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
+The bound wave functions associated with the Morse potential are employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
 $\mu_{ij}$ $=\int\psi^{*}_{i}(r)\mu(r)\psi_{j}(r)\mathrm{d}r,$ (7)
 where $\psi_{i}(r)$ represents the $i$th vibrational level wave function, and the dipole function $\mu(r)$ models the nonlinear change in the (electronic ground-state) molecular electrical dipole with the bond length $r$ [52].
-We investigated the IR radiative dissociation of hydrogen fluoride, HF, and lithium hydride, LiH.
-The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-4 were obtained from previous work on laser-induced HF dissociation[41].
-The LiH dipole function was interpolated from ab initio quantum chemical computations (CCSD(T)/aug-cc-pVTZ, see Supplementary Information for further details).
-Morse potential parameters include the equilibrium bond length ($r_{e}$), electronic ground-state dissociation energy ($D_{e}$), and harmonic frequency ($\omega_{e}$) of each molecule and are listed in Table 1.
-Table 1: Parameters used for Morse potential of HF [42, 43, 44, 41] and LiH [45].
-Constant HF LiH
-$D_{e}$ (eV) 6.123 2.641
-$r_{e}$ (Å) 0.926 1.595
-$\nu_{e}$ (cm-1) 4138 1405
-$\nu_{e}\chi_{e}$ (cm-1) 86.7 28.3
-$m_{r}$ (a.u) 0.95 0.88
-The examined dissociation processes (via the most energetic bound state) have reduced room temperature rates relative to typical BIRD experiments involving polyatomic ions [27, 46, 29, 22] due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation proceeding via the highest energy level.
-This feature motivated our choice of examining diatomic BIRD at temperatures $T=\mathcal{O}(\hbar\omega_{e}/k_{\text{B}})$, an order of magnitude greater than room temperature.
-The SI shows that the trends discussed below hold for a wide interval of temperatures.
-Despite their different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior.
+The rates $k_{\text{loss}}^{i}$ for radiative absorption into the continuum by each bound state $i$ are obtained by a simple generalization of Eqs. 5 as discussed in the Supplementary Material.
+We investigated the thermal IR radiative dissociation of the NaLi molecule.
+The electrical dipole function and Morse potential were obtained via interpolation from previous CCSDT/cc-pCVQZ calculations available in the literature [53] (see Supplementary Information for further details).
+Morse potential parameters are listed in Table 1.
+Table 1: Parameters characterizing the Morse potential associated to NaLi.
+Parameter Notation (unit) Value
+Electronic ground-state dissociation energy $D_{e}$ (eV) 0.882
+Equilibrium bond length $r_{e}$ (Å) 2.895
+Harmonic frequency $\omega_{e}~{}\left(\text{cm}^{-1}\right)$ 257.4
+Anharmonicity $\omega_{e}\chi_{e}~{}\left(\text{cm}^{-1}\right)$ 2.33
+Reduced mass $m_{r}$ (a.u) 5.331
 The transition rates $k_{ij}$ are proportional to the electromagnetic density of states $D(\omega)$, suggesting BIRD rates can be controlled by placing the reactive molecule inside a microcavity.
 To examine this scenario, we adapt the free space formalism described above to arbitrary photonic devices by replacing $D_{0}(\omega)$ in Eqs.
-3 and 5 with the appropriate DOS for a weakly coupled microcavity $D_{\text{C}}(\omega)$.
+4 and 6 with the appropriate DOS for a weakly coupled microcavity $D_{\text{C}}(\omega)$.
 In a planar empty Fabry-Perot microcavity composed of two parallel mirrors (with unit reflectivity for simplicity) separated by a distance $L_{C}$, [Fig.
@@
 where $\lfloor x\rfloor$ denotes the floor function.
+This work focuses on establishing an upper limit for microcavity effects on BIRD by assuming perfectly reflecting mirrors.
+Real microcavities are leaky, so to assess the robustness of our predictions, we also computed the photon density of states and corresponding quantum state transition rates in imperfect microcavities with finite reflectivity and absorbing metallic mirrors using the electromagnetic field dyadic Green function [56] (see Supplementary Information for details).
 Lastly, we considered the BIRD scenario in which the diatomic interacts weakly with the electromagnetic component of polariton modes emergent from vibrational strong coupling between the microcavity modes and a material system (solid-state or molecular ensemble) with sufficiently large IR oscillator strength.
-We emphasize this material system does not undergo BIRD; its role is to generate polariton normal-modes with a tunable density of states and thermal energy density, as shown in Figs.
+We emphasize that the polariton modes arise from collective strong coupling between the cavity and a host material with substantial infrared oscillator strength.
+The reactive molecule experiences modified radiative dynamics due to changes in the photon density of states induced by the formation of polaritons, as illustrated in Figs.
 1(d) and (e).
+Consequently, in this study, the interaction strength between our reactive diatomic and the radiation field is obtained from first principles without any scaling parameters.
+While this polariton-assisted scenario poses challenges for experimental realization as BIRD generally requires low-density, collision-free conditions, it is introduced here as a theoretical construct that isolates the effect of polariton-modified photonic environments on molecular infrared radiative dissociation.
+Nonetheless, one could envision suppressing collisions by confining reactive molecules using transparent partitions or external fields, while maintaining exposure to the confined electromagnetic modes[41, 42, 43].
+A detailed treatment of these experimental considerations, as well as the role of direct interactions between host and reactive molecules, lies beyond the scope of the present work and warrants future investigation.
 The formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD (see SI) by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
@@
 […]
-$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-\frac{m\pi c}{L_{C}}],$ (11)
-where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$
+$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-m\pi c/L_{C}],$ (12)
+where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the photon content of the polariton modes with frequency $\omega$
 $P_{C}(\omega)=\frac{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}}{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}+\omega^{2}_{C}(\omega)\Omega_{R}^{2}},$ (13)
@@
 As previous work shows [70], introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$.
-Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on the type of disorder and coherence loss mechanism.
-This motivates our decision to use the formally exact expression presented in Eq.
+Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on disorder and coherence loss mechanisms.
+Similarly, the stopgap is known to contain weakly coupled states in disordered systems causing $D_{P}(\omega)$ to become nonzero in this region.
+However, the photon density of states remains lower in the stopgap in comparison to the weak coupling regime.
+The robustness of the enhanced and reduced photon-weighted polariton density of states at frequencies $\omega$ approaching $\omega_{M}$ from the left and right, respectively, support our choice to to use the formally exact expression presented in Eq.
 12 and interpret our results as upper bounds on the effect of polaritons on BIRD.
+The relevance of these bounds to experiments will be shown to depend on the collective light-matter interaction strength and the detuning between transitions of interest and $\omega_{M}$ as discussed in detail in Results and Discussion.
 III   Results and Discussion
-Microcavity-assisted blackbody infrared radiative dissociation
+III.1      Microcavity-assisted blackbody infrared radiative dissociation via doorway state
 Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity.
-The tunable electromagnetic environment afforded by the microcavity allows BIRD rates to be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the modified photon DOS.
+The tunable electromagnetic environment afforded by the microcavity allows BIRD rates to be either enhanced or suppressed relative to the free space, depending on which transitions are affected by the modified photon DOS.
 In Fig.
-2, BIRD rates in the weak coupling regime are shown relative to the free space rate ($k_{0}=2.70\times 10^{-6}$ s-1 for HF at $T/\omega_{e}=0.67$ corresponding to $T=4000~{}\text{K}$ and $k_{0}=1.64\times 10^{-5}$ s-1 for LiH at $T/\omega_{e}=0.99$ corresponding to $T=2000~{}\text{K}$).
-For small microcavity lengths, i.e., $L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH, a mild but consistent enhancement of BIRD rates is observed.
+2(a), BIRD rates in the weak coupling regime are shown relative to the free space rate at $T=400~{}\text{K}$.
+For small microcavity lengths $L_{C}<25$ $\mu$m, a mild but consistent enhancement of BIRD rates is observed.
 At these lengths, an overall increase in the microcavity electromagnetic DOS over a wide frequency range covering several important transitions is associated with the observed enhancement.
-The ratios of microcavity BIRD rates to free space BIRD rates quickly approach 1.0 as $L_{C}$ increases beyond a few microns.
+The ratios of microcavity BIRD rates to free space BIRD rates approach 1 as $L_{C}$ increases beyond a few tens of microns.
 This is expected since the microcavity DOS converges to the free space limit when $L_{C}\rightarrow\infty$.
-The complex oscillatory pattern in the BIRD rates relative to free space shown in Fig.
-2 arises from the interplay of discrete vibrational transitions enhanced or suppressed depending on $L_{C}$.
+On top of the main feature described above, we also observe an oscillatory pattern in the BIRD rates relative to free space shown in Fig.
+2(a).
+This pattern arises from the interplay of discrete vibrational transitions enhanced or suppressed depending on $L_{C}$.
 As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
-Figure 2: Ratio of BIRD rate inside a microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for HF and LiH molecules.
-Red circles highlight specific lengths further discussed in Fig.
-3.
+Figure 2: (a) Ratio of BIRD rate inside a nearly empty microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for NaLi.
+Red circles highlight specific lengths further discussed in (b) and (c).
+Microcavity photon DOS normalized to the free space DOS for $L_{C}=26.3$ and $26.8$ $\mu$m, in (b) and (c), respectively.
+The dotted vertical lines indicate the location of the molecular transition frequencies corresponding to the most relevant overtone transitions, with color coding to distinguish between suppression (red) and enhancement (green).
 We conducted a sensitivity analysis of the dissociation rate to interpret the origin of the oscillatory behavior in Fig.
-2 and identify the most consequential radiative transitions involved in the BIRD of HF and LiH.
-This analysis measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details).
-For HF dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 23$, with $i\in\{16,17,19\}$ (in the relevant 1500 -- 5000 cm-1 frequency range) are the most consequential for the BIRD rate.
-Similarly, for LiH, the relevant transitions are $i\rightarrow 29$, with $i\in\{21,22,23,25\}$.
-In both cases, the identified overtones take the vibrational mode to the highest energy-bound state (23 for HF and 29 for LiH), from which dissociation rapidly occurs.
+2(a) and identify the most consequential radiative transitions involved in the BIRD of NaLi.
+This analysis measures how impactful an enhancement in the photon DOS at a specific transition frequency is on the dissociation process (see Supporting Information for more details).
+For NaLi dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 54$, with $i\in\{42,43,48\}$ are the most consequential for the BIRD rate.
+The identified overtones take the vibrational mode to the highest energy-bound state, from which dissociation rapidly occurs.
 We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths.
 See Supporting Information for a brief discussion of results obtained without overtone transitions.
-Figure 3: Microcavity photon DOS normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig.
+Equipped with knowledge of the most relevant transitions for the investigated BIRD of NaLi via the most energetic bound state, we can explain the microscopic origin of the intricate pattern in Fig.
+2(a) considering how the microcavity DOS changes around the previously mentioned overtones.
+Figs.
+2(b) and 2(c) present the ratio of microcavity DOS to the free space for selected values of $L_{C}$ where suppression or enhancement of BIRD was verified in Fig.
+2(a) (red circles).
+Overtone frequencies are highlighted with dotted vertical lines, color coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity.
+In Fig.
+2(b) ($L_{C}=26.3$ $\mu$m), the most relevant NaLi overtones are suppressed, and this leads to the local minimum observed in $k_{c}/k_{0}$ in Fig.
 2.
-The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color coding to distinguish between suppression (red) and enhancement (green).
-Panel (a) shows that the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced relative DOS at the highlighted overtones.
-Panel (b) shows the mild enhancement of HF BIRD in a microcavity with $L_{C}=3.21~{}\mu\text{m}$ can be ascribed to the greater microcavity photon DOS at the frequencies of the highlighted overtones.
-Panels (c) and (d) show the analogous scenario for LiH, where BIRD is observed to be suppressed at $L_{C}=9.4~{}\mu\text{m}$ but enhanced at $L_{C}=9.6~{}\mu\text{m}$, primarily due to the variation of the microcavity photon density of states at the selected $i\rightarrow 29$ overtone frequencies.
-Equipped with the knowledge of the most relevant transitions for the investigated BIRD of HF and LiH via the most energetic bound state, we can readily explain the microscopic origin of the intricate pattern in Fig.
-2 by considering how the microcavity DOS changes around the previously mentioned overtones.
-Fig.
-3 presents the ratio of the microcavity DOS to the free space for selected values of $L_{C}$ where BIRD suppression or enhancement was verified in Fig.
-2 (red circles).
-Overtone frequencies are highlighted with dotted vertical lines, color-coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity.
-In Fig.
-3(a) ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig.
-2.
-In contrast, when the microcavity length slightly increases to $L_{C}=3.21$ $\mu$m [Fig.
-3(b)] the photon DOS is enhanced at the $16\rightarrow 23$ and $19\rightarrow 23$ transitions frequencies, and we observe a corresponding leap in the BIRD rate in Fig.
-2.
-Similarly, in Fig.
-3(c) ($L_{C}=9.4~{}\mu\text{m})$, all of the relevant LiH overtones but $21\rightarrow 29$ are suppressed, leading to a reduced BIRD rate in this microcavity relative to free space, whereas Fig.
-3(d) shows that at $L_{C}=9.6~{}\mu\text{m}$ most of the relevant LiH overtones are enhanced, thus explaining the sudden increase in BIRD rate at $L_{C}=9.6~{}\mu\text{m}$ relative to $L_{C}=9.4~{}\mu\text{m}$.
+In contrast, when the microcavity length slightly increases to $L_{C}=26.8$ $\mu$m [Fig.
+2(c)] the photon DOS is enhanced at the $42\rightarrow 54$ transition frequency, and we observe a corresponding leap in the BIRD rate in Fig.
+2(a).
 Introducing a finite linewidth to the considered radiative transitions would smooth out the oscillations observed in Fig.
-2 and dampen their amplitude.
-Consequently, the weakly coupled microcavity effect on BIRD would be significant only at lengths less than a few microns.
+2(a) and dampen their amplitude.
+Consequently, the weakly coupled microcavity effect on BIRD would be significant only at lengths less than a few tens of microns.
 Note that suppression effects are generally minor in Fig.
-2 because the dissociative process has many pathways.
+2(a) because the dissociative process has many pathways.
 Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always negligible as the molecular system explores alternative dissociation pathways.
-Our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
-Polariton-assisted blackbody infrared radiative dissociation
+Our analysis of BIRD in weak coupling reveals two main results that we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
+We conclude this section by noting that BIRD simulations with lossy mirrors (SI Sec.
+8) show (i) broadening and suppression of the sharp features in Fig.
+2 and (ii) enhanced dissociation rates relative to Fig.
+2(a) at short $L_{C}$ due to the presence of evanescent modes with significant amplitude near the imperfect metal interfaces[7, 61].
+Overall, lossy mirrors suppress the oscillatory behavior of the relative BIRD rate with microcavity length while preserving the order-of-magnitude effects predicted for perfect mirrors.
+This confirms that, modulo the fast oscillations in Fig.
+2(a), the results presented in this Section are representative of experimentally accessible photon resonators.
+III.2      Polariton-assisted blackbody infrared radiative dissociation via doorway state
 To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes.
@@
 1 (e)] and the electromagnetic DOS.
-We explore the implications for diatomic BIRD via the most energetic bound state below.
+Importantly, the reactive NaLi molecule remains weakly coupled to the confined electromagnetic field.
+We explore the implications for diatomic BIRD via the most energetic bound state in the following.
 Figure 3: Ratio of polariton-assisted BIRD rates to free space BIRD rates.
 In the polaritonic case, the diatomic molecule is embedded in a strongly coupled microcavity with variable host material with frequency $\omega_{M}$.
-The top panels, (a) and (b), correspond to HF, and the bottom panels, (c) and (d), show results for LiH.
-In (a) and (c), the collective light-matter interaction strength, $\Omega_{R}$, is fixed at 400 and 200 cm-1, respectively.
+In (a), the collective light-matter interaction strength, $\Omega_{R}$, is fixed at 200 cm-1.
 The horizontal dashed gray and green lines indicate where the dissociation rates equal $k_{0}$ and $k_{c}$, respectively.
-The insets in (a) and (c) show zoomed-in views around the overtones $17\rightarrow 23$ for HF and $22\rightarrow 29$ for LiH.
-In (b) and (d), relative BIRD rates are shown for different Rabi frequencies ($\Omega_{R}$), depicting the dependency on the detuning between the host molecule and the diatomic transition energies.
-Fig.
-4 presents the ratio of polariton-assisted BIRD to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$.
-These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 by fixing $L_{C}$ at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH respectively.
-As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime.
-We observe prominent peaks for HF and LiH where the relative dissociation rate is notably enhanced.
-Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$.
-Here, we focus on the case where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH.
-In contrast, the inset of Figs.
-4(a) and (c) show that when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed.
-This suppression occurs because as $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the overtone transition falls into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig.
-1e).
-However, the BIRD suppression is relatively small since, as discussed earlier, the reaction may still proceed through alternative pathways.
-The effect of the collective coupling strength ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig.
-4(b) and (d).
-When $\omega_{M}-\omega_{i\rightarrow j}$ is around 0.01-1 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates.
-However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively.
-This plateau appears to be insensitive to the magnitude of the Rabi splitting.
-It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig.
-5.
-Figure 5: Photon-weighted polariton density of states $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with length $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at frequency $\omega_{M}=1401$ cm-1.
+The minimum difference between $\omega_{M}$ and $\omega_{i\rightarrow j}$ used here is 1 cm-1 to match panel (b).
+The inset in (a) shows a zoomed-in view around the overtone $42\rightarrow 54$.
+In (b), relative BIRD rates are shown for different Rabi frequencies ($\Omega_{R}$), depicting the variation of the relative rate enhancement with the detuning between the host molecule and the diatomic transition energies.
+Figure 4: Photon-weighted polariton density of states $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with length $L_{C}=26.8~{}\mu$m strongly coupled to a material with a bright transition at frequency $\omega_{M}=380$ cm-1.
 The curves show the effect of different Rabi splitting values ($\Omega_{R}$) on the photon-weighted polariton DOS, which is relevant for radiative processes mediated by polaritonic systems.
@@
 This trend explains the impact of the collective light-matter interaction strength on the observed polariton-assisted bond infrared dissociation (BIRD) enhancement shown in Fig.
-4.
+3.
+Fig.
+3 presents the ratio of polariton-assisted BIRD to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$ and fixed microcavity length $L_{C}$ at 26.8 $\mu\text{m}$ where we observed enhanced rates.
 In Fig.
-5, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=9.6$ $\mu$m and $\omega_{M}=v_{22\rightarrow 29}(\text{LiH})=1401$ cm-1.
-The case where $\Omega_{R}=0$ represents the weak coupling limit and is included here to directly compare $D_{P}(\omega)$ with $D_{C}(\omega)$.
+3(a), we observe prominent peaks where the relative rates are notably enhanced.
+These peaks occur when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to the specific overtone transition frequencies ($\omega_{i\rightarrow j}$).
+The asymmetry in the relative dissociation rate profile around the observed maxima, shown in more detail in the inset of Fig.
+3(a), emerges as a result of the formation of the polariton stopgap region.
+When $\omega_{M}$ is close but smaller than $\omega_{i\rightarrow j}$, this transition falls within the stopgap region and is completely suppressed as the density of states becomes zero at that frequency.
+Nonetheless, as discussed for the weak coupling regime, a strong suppression of reaction rates is not observed due to the many alternative pathways the reaction can access.
+Focusing on the $42\rightarrow 54$ overtone ($\omega_{42\rightarrow 54}=380~{}\text{cm}^{-1}$) which is the most consequential transition for the reaction, the effect of a variation in $\omega_{M}-\omega_{42\rightarrow 54}$ at various collective coupling strengths ($\Omega_{R}$) on BIRD rates is examined in Fig.
+3(b).
+Increasing $\Omega_{R}$ leads to a significant enhancement of the polariton-assisted BIRD rates, reaching a factor of 40 at the large Rabi frequency $\Omega_{R}=$200 cm-1 and small $\omega_{M}-\omega_{42\rightarrow 54}$.
+As this frequency difference increases, the enhancement drops: at 3 cm-1, the enhancement is about 12-fold; at 5 cm-1 detuning, around 7-fold; and beyond 7 cm-1, the enhancement persists around 4-fold.
+This reduction in dissociation rate enhancement with increasing $\omega_{M}-\omega_{42\rightarrow 54}$ follows from the behavior of the photon-weighted polariton DOS near the stopgap as we discuss next.
+In Fig.
+4, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=26.8$ $\mu$m and $\omega_{M}=\omega_{42\rightarrow 54}=380$ cm-1.
+The case where $\Omega_{R}=0$, representing the weak coupling limit, is included here for direct comparison with $D_{C}(\omega)$.
 It can be seen from Fig.
-5 that the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap.
-Most remaining features, e.g., the oscillations in $D_{P}(\omega)$ with $\omega$, are also observed in weak coupling and can be understood from Eq.
-8.
-The various curves in Fig.
-5 illustrate how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ expand as $\Omega_{R}$ increases.
-As a result, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting.
-The singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate.
-However, our results demonstrate that this singularity does not lead to arbitrarily fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting.
-Therefore, the BIRD rates reported as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here.
-In experiments, imperfections in the electromagnetic device and the strongly coupled material smooth out the photon-weighted DOS near the stopgap.
-Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder [63].
-The introduction of a UV cutoff, loss, and dephasing pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig.
-4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc).
-Including a finite linewidth for the diatomic vibrational transitions relaxes the strict resonance requirement discussed above ($\omega_{M}\rightarrow\omega_{i\rightarrow j}$), and the overall enhancement will depend on the spectral overlap between $\omega_{i\rightarrow j}$ and $D_{P}(\omega)$.
-Since the molecular linewidth and cavity imperfections are expected to be limiting factors, it is clear that our reported results represent theoretical upper bonds for polariton-assisted BIRD rates for the reaction mechanism investigated here.
-Before ending, we emphasize our observation of the same qualitative behavior for polariton-mediated BIRD in LiH and HF.
+4 that the photon-weighted polariton DOS is only significantly different from the bare microcavity DOS in a frequency range around $\omega_{M}$ defined as $\{\omega_{M}-\alpha(\Omega_{R})<\omega<\omega_{M}+\alpha(\Omega_{R})\}$, where $\alpha(\Omega_{R})$ is a positive frequency that depends on $\Omega_{R}$ and satisfies $D_{P}[\omega_{M}\pm\alpha(\Omega_{R})]\approx D_{C}[\omega_{M}\pm\alpha(\Omega_{R})]$.
+Outside this range, $D_{P}(\omega)$ retains features of the weak coupling case, such as oscillations in $\omega$, which can be understood from Eq.
+9.
+Fig.
+4 further illustrates how increasing $\Omega_{R}$ broadens the stopgap region and expands $\alpha(\Omega_{R})$, leading to a wider frequency range for which $D_{P}(\omega)\gg D_{C}(\omega)$.
+As a result, depending on the Rabi splitting, strong coupling can induce changes in the rates of multiple vibrational transitions.
+The singular behavior of $D_{P}(\omega)$ at $\omega_{M}$ is originated by the implicit assumption of our model that off-resonant photons with arbitrarily large energy mix with the strongly coupled matter system excitations and form polaritons.
+These interactions, albeit weak, lead to the formation of lower polariton (LP) modes with very small photon contents and frequencies arbitrarily close to $\omega_{M}$.
+This behavior is reflected in the expression for the photon-weighted polariton DOS in Eq.
+12, where the photon content appears in the numerator and the polariton group velocity in the denominator.
+As the frequency $\omega$ approaches $\omega_{M}$, the group velocity vanishes faster than the photon content, resulting in an arbitrarily large photon DOS.
+This divergence in the DOS manifests itself as an accumulation of LP states near $\omega_{M}$, driven by the off-resonant interaction between matter and the microcavity radiation field.
+In other words, as the in-plane wave vector $k$ increases, the LP transition frequencies approach $\omega_{M}$, yet they never become exactly equal in our model due to the presence of the stopgap.
+Consequently, within any arbitrarily small frequency interval $\omega_{M}-\delta<\omega<\omega_{M},~{}~{}\delta>0$, a large number of LP states exist, leading to a singularity in the photon-weighted polariton DOS at $\omega_{M}$.
+The off-resonant light-matter interactions responsible for the singularity of $D_{P}(\omega)$ at $\omega_{M}$ can be eliminated by imposing an effective frequency cutoff for polaritons formed from high-energy microcavity modes by setting a minimum nonvanishing $\delta=\omega_{M}-\omega_{42\rightarrow 54}$.
+For example, when $\delta=10$ cm-1 photon modes across 6 branches with $\omega_{C}(m,q)=945$ cm-1 contribute to $D_{P}(\omega)$, while when $\delta=1$ cm-1, $D_{P}(\omega)$ contains contributions from 15 microcavity branches with $\omega_{C}(m,q)=2780$ cm-1.
+Therefore, by choosing a nonvanishing threshold for $\delta$ we also set a maximum energy (cutoff) for radiation states that contribute to the photon-weighted polariton DOS.
+Fig.
+3 shows the removal of an increasing number of polaritons formed from highly off-resonant photons reduces the relative BIRD rate enhancement to an extent that depends on the Rabi frequency and the minimum $\delta$.
+In experiments, imperfections in the strongly coupled material system and in the electromagnetic device (SI, Sec.
+8) smooth the photon-weighted DOS near the stopgap.
+Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet (UV) cutoff and material disorder [70].
+The introduction of a UV cutoff, cavity loss, and dephasing pathways for the matter polarization field a̱re therefore expected to lead to reduced BIRD enhancements relative to those in Fig.
+3 depending in a non-universal way on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc).
+Still, given the modest impact of leaky mirrors on the order of magnitude of the microcavity effect on the BIRD rates discussed in Sec.
+III.A and SI Sec.
+8, we expect the order of magnitude of polariton-assisted BIRD enhancement to be similarly robust to the introduction of microcavity losses characteristic of experimentally accessible devices.
+Note the singularity in the photon-weighted polariton DOS at $\omega_{M}$ enables a particular radiative transition to become much faster than the corresponding free space rate.
+However, it does not lead to arbitrarily fast reaction rates, since the complex multistep reactive process ensures that other transitions quickly become rate-limiting (see the Supporting Information for a saturation analysis).
+Therefore, the reported BIRD rates obtained when $\omega_{M}$ approaches selected $\omega_{i\rightarrow j}$ (Fig.
+3) provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here.
+Before ending we emphasize our observation of the same qualitative behavior for polariton-mediated BIRD in other diatomic species (see SI).
 Since these chemical species have significantly different energetics, we anticipate the main trends reported in this work also apply to other diatomic and polyatomic molecules, particularly in cases where dissociation occurs through a specific doorway state.
+III.3      Dissociation via radiative bound-continuum transitions
+In the previous sections, we examined BIRD mediated by the most energetic bound-state.
+Gas-phase dissociation may also be promoted by direct decay from any bound-state $i\leq i_{\text{max}}$ into the dissociative continuum.
+This scenario is examined in this section, where we generalize our previous treatment to include direct (irreversible) dissociation pathways for each bound-state $i$ with rate $k^{i}_{\text{loss}}$ as described in Methods and Supporting Information.
+Fig.
+5(a) shows the relative dissociation rate, $k_{C}/k_{0}$, for a diatomic molecule inside a nearly empty microcavity as a function of its length.
+In contrast to the doorway mechanism discussed in Sec.
+IIIA, which exhibits pronounced oscillations, the behavior here is noticeably simpler, yet may be understood with the same tools previously employed to rationalize the observed trends in Fig.
+2.
+For example, from the normalized photon DOS at microcavity lengths $L=10.5~{}\mu\text{m}$ and $L=15.5$ $\mu$m presented in Figs.
+5(b,c) corresponding to local minimum and maximum in Fig.
+5(a), respectively, we observe the ten most influential transitions, identified through sensitivity analysis, lie between 330–420 cm-1.
+At $L=10.5~{}\mu\text{m}$, these transitions fall within a frequency range where the DOS is suppressed relative to free space, whereas at $L=15.5~{}\mu\text{m}$, they match with a region of enhanced DOS.
+This variation accounts for the observed changes and local extrema in the microcavity dissociation rates in Fig.
+5(a).
+Figure 5: (a) Ratio of BIRD rate inside a nearly empty microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for NaLi.
+Microcavity photon DOS normalized to the free space DOS for $L_{C}=10.5$ and $15.5$ $\mu$m, in (b) and (c), respectively.
+The dotted vertical lines indicate the location of the molecular transition frequencies corresponding to the most relevant overtones, with color coding to distinguish between suppression (red) and enhancement (green).
+Note when the diatomic dissociation occurs exclusively via the doorway state (as in Sec.
+III.A.), a few overtone transitions transferring population to the highest energy bound state control the reaction rate.
+However, when dissociation can proceed directly from each bound state, sensitivity analysis shows that numerous $\Delta v=+2$ transitions (e.g., $12\rightarrow 14$, $13\rightarrow 15$, and $14\rightarrow 16$) can have a significant impact in the reaction rate.
+This feature challenges the assignment of dominant reaction pathways, yet relaxes the conditions for enhancement or minor suppression of BIRD in a weakly coupled microcavity.
+Fig.
+6(a) presents the relative polariton-mediated BIRD rate, $k_{P}/k_{0}$, as a function of the host material frequency $\omega_{M}$ under strong coupling conditions with a collective light-matter interaction strength of $\Omega_{R}$ = 125 cm-1 .
+Here, $L_{C}$ is fixed at 5 $\mu$m, and decay contributions from all vibrational states are included (unlike the results in Sec.
+III.B which considered only decay from the highest bound state).
+In contrast to the isolated sharp peaks observed when the reaction proceeds via the loosest bound state (Fig.
+3), the introduction of multiple direct dissociative channels via bound-continuum transitions leads to a broader distribution of rate enhancements over a wide range of $\omega_{M}$.
+While resonant features persist in the reaction rate profile corresponding to particular overtone transitions (Fig.
+3a), the observed enhancements largely reflect cumulative contributions of decay channels proceeding directly from many intermediate vibrational bound states.
+To estimate the effect of broad vibrational linewidths on the relative polariton-assisted BIRD rates, we applied locally estimated scatterplot smoothing (LOESS)[71, 72, 73], a generalized moving average method.
+The resulting LOESS curve, shown as a solid line in Fig.
+6, reveals the underlying polaritonic enhancement of the BIRD rates while smoothing over sharp features that arise from approximating the continuum states of the Morse potential using a dense but discrete set of box states.
+This averaging is also supported by the presence of physical broadening mechanisms, including thermal rotational population and weak rovibrational interactions, which naturally contribute to infrared linewidths.
+The smoothed polariton-assisted BIRD rates reveal that at $\Omega_{R}$ = 125 cm-1, enhancement is possible when the host material frequency falls within the 300–500 cm-1 range.
+A mild suppression compared to the weak coupling case ($k_{P}/k_{C}<1$) is also possible when the host material frequency falls within the 200–300 cm-1 range.
+We attribute these features to the modification of bound-continuum transition rates ($k_{\text{loss}}^{i}$) as demonstrated in Fig.
+6(b) which shows these for bound vibrational levels near the dissociation threshold.
+When $\omega_{M}=230~{}\text{cm}^{-1}$, the decay rates for intermediate bound states are slightly suppressed compared to weak coupling.
+In contrast, when $\omega_{M}=400~{}\text{cm}^{-1}$, decay rates for states $i=43$–47 are enhanced due to the modified photon weighted polariton DOS.
+Therefore, in Fig.
+6(a), the minimum near $\omega_{M}=230~{}\text{cm}^{-1}$ arises from suppressed decay channels due to a locally depleted polariton DOS, while the maximum near $\omega_{M}=400~{}\text{cm}^{-1}$ results from enhanced decay rates promoted by enhanced polariton DOS.
+When the reaction is allowed to proceed through this mechanism, which is sensible for gas phase systems, maximum enhancements are reduced compared to the BIRD mediated by the highest energy bound state (or any particular specific doorway state).
+However, demanding resonance conditions are also lifted affording a more distributed enhancement profile that may offer a practical advantage for experimental realizations.
+Figure 6: (a) Ratio of polariton-assisted BIRD rates to free-space BIRD rates for $\Omega_{R}=125~{}\text{cm}^{-1}$.
+The horizontal dashed green line corresponds to the weakly coupled microcavity relative rate $k_{C}/k_{0}$.
+The solid line was computed using a LOESS moving average[71, 72, 73], which highlights the overall trend while smoothing out sharp resonance features.
+(b) Decay rates from bound vibrational states ($i=40$–50) to the continuum under weak coupling ($L_{C}=5~{}\mu\text{m}$), and under strong coupling for different $\omega_{M}$ values in the polaritonic environment.
 IV   Conclusions
@@
 In microcavities where $L_{C}\leq 25~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates.
-As $L_{C}$ increases beyond a few microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones.
+As $L_{C}$ increases beyond a few tens of microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones.
 Further increases in $L_{C}$ causes the BIRD rates to approach the free-space limit.
 In the strong coupling regime, where polaritons mediate BIRD, we reported $O(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process.
-This enhancement is due to the emergence of a singularity in the photon-weighted polariton DOS, which massively increases transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below.
-The singularity is attributed to the lack of a high-energy cutoff for the electromagnetic field and the lossless character of the strongly coupled system.
-Nevertheless, while the singular behavior of the photon-weighted polariton DOS is absent from experiments, we have argued its effects are expected to persist to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder.
-Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced diatomic thermal radiative dissociation changes via the most energetic bound state.
-Our work focused on how optical microcavities can modify BIRD reactions of diatomic molecules via a particular mechanism.
-Despite focusing on these specific systems, the main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly.
-We expect the analysis introduced here will be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.
+This enhancement is due to the emergence of a Rabi frequency dependent enhanced photon-weighted polariton DOS at a targeted frequency, which enables the increase of transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below.
+The largest enhancements are obtained when photon modes highly off resonant with the host molecular system are assumed to form polaritons, yet as we showed, sizable polariton-assisted BIRD rate enhancements remain achievable when these highly off resonant interactions are cutoff to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder.
+Ultimately, the BIRD enhancement factors here reported must be viewed as upper bounds for selected polariton-induced diatomic thermal radiative dissociation via the most energetic bound state.
+The inclusion of bound-to-continuum transitions in our analysis alters the dissociation dynamics compared to the case where the reaction proceeds via a doorway state (the highest energy bound state).
+Dissociation via radiative bound-continuum transitions reveals a broader and smoother enhancement profile under microcavity length variation in the weak coupling regime.
+While the magnitude of the enhancement remains similar, the key difference lies in the involvement of a much greater number of vibrational transitions relative to the doorway mechanism.
+In the polariton-assisted case, relative rate enhancement is generally smaller, but it no longer requires precisely matching a particular diatomic transition frequency to the host (strongly coupled) material.
+Instead, the multiplicity of direct dissociation pathways makes polariton-enhanced BIRD inherently more robust, enabling broadband enhancement across a wide spectrum of host material frequencies $\omega_{M}$.
+Our work focused on how optical microcavities can modify diatomic blackbody infrared radiative dissociation.
+While our analysis ignores collisional energy transfer, and therefore may be viewed as providing upper bounds to the microcavity effect on diatomic BIRD rates, the main qualitative conclusions expressed above are expected to apply broadly to systems where thermal radiation acts as a relevant energy source.
+We anticipate that the framework developed here will aid in understanding and controlling thermal radiative-induced dissociation processes in more complex, polyatomic systems.
+Supporting Information. Results for the HF and LiH molecules, temperature analysis, and additional computational details, including derivations, model parameters, and sensitivity analysis.
 Acknowledgments.